# Rolled Gaussian process models for curves on manifolds

### v1

@@ Abstract
Given a planar curve, imagine rolling a sphere along that curve without slipping or twisting, and by this means tracing out a curve on the sphere. It is well known that such a rolling operation induces a local isometry between the sphere and the plane so that the two curves uniquely determine each other, and moreover, the operation extends to a general class of manifolds in any dimension. We use rolling to construct an analogue of a Gaussian process on a manifold starting from a Euclidean Gaussian process. The resulting model is generative, and is amenable to statistical inference given data as curves on a manifold. We illustrate with examples on the unit sphere, symmetric positive-definite matrices, and with a robotics application involving 3D orientations.
@@ 1 Introduction
1.1 Background and motivation
Many modern applications generate data where the statistical unit is a curve on a manifold, $M$. An interesting example is robotics: in a given task, a robot arm’s end-effector traces a curve in a space of orientations. We consider one such example later, with data as smooth curves $t\mapsto x(t)\in M$, where $M$ is the manifold of rotation matrices. Despite prevalence of such data (e.g. Su et al., 2014; Zhang et al., 2018), there is a dearth of suitable generative and statistical models that are simple and tractable.
@@ 1.1 Background and motivation
The four maps are compatible with the intrinsic geometry of $M$, do not introduce any distortions, and are optimal in a sense clarified below. This is in contrast to identifying $\mathbb{R}^{d}$ with a tangent space at a single fixed point on $M$, and using tangent space projections between the tangent space and $M$, and vice versa, which causes distortions. Our maps provide a natural way to define a covariance, and hence a Gaussian process model, on $M$, starting with one in $\mathbb{R}^{d}$.
1.2 Related works, and contributions
@@
“Rolling” of a curve in $\mathbb{R}^{d}$ onto $M$, and its reverse operation, “unrolling”, originate from mathematical geometry, where they are known as antidevelopment and development respectively; see Kobayashi and Nomizu (1996) and Sharpe (2000). Jupp and Kent (1987) first exploited the idea in statistics, to define smoothing curves for time-indexed data points on the unit sphere. Kume et al. (2007) adapted their technique to shape spaces, and later Kim et al. (2021) generalized to Riemannian manifolds. These works, however, did not consider statistical models for inference.
@@
The notion of stochastic development has been used to define generative diffusion models for curves on $M$ (e.g. Sommer, 2019, 2020) based on stochastic differential equations in $\mathbb{R}^{d}$. Markovian property of the models do not allow for rich covariance structures, and computational challenges abound: exact simulation is infeasible, numerical approximations are costly, and likelihood functions lack closed-form expressions, complicating inference.
Notions of covariance of a random curve $x$ on $M$ with respect to a fixed curve $\gamma$ have been considered. The extrinsic approach by Dai and Müller (2019) for embedded submanifolds $M$ results in a covariance that depends on the embedding. To define an intrinsic covariance operator of a random vector field along $\gamma$, Lin and Yao (2019) used the tensor Hilbert space of tangent spaces along $\gamma$, while Shao et al. (2022) used an abstract notion of a covariance bundle. The common focus of these papers is on principal component analysis with $\gamma$ as the Fréchet mean curve, albeit under different sampling regimes, and models for inference were not considered. Methodology for curves with values in a general metric space that do not explicitly use the differentiable structure of the manifold $M$ have been developed (Dubey and Müller, 2020, 2022).
Mallasto and Feragen (2018) defined a model for $x$ on $M$ by wrapping, under the exponential map, a Gaussian vector field along a mean curve $\gamma$ with covariance as in Lin and Yao (2019). A similar approach was adopted recently by Liu et al. (2024) and Wang et al. (2024). The issue with using such models for inference is two-fold: as with $\mathbb{R}^{d}$-valued Gaussian processes, the mean curve $\gamma$ and the covariance have to be jointly estimated; however, the strategies used in $\mathbb{R}^{d}$, such as using a basis representation for $x$, or a parametric covariance structure, are difficult to devise. In the above works, the mean function is either pre-specified or estimated separately.
The main contributions of this paper are: (i) to use rolling and wrapping with respect to coordinate frames to specify how a Gaussian process model in $\mathbb{R}^{d}$ can be identified with a “rolled Gaussian process” model on $M$, while establishing its invariances and equivariances to choice of frames; (ii) to establish conditions for the Fréchet mean of the rolled Gaussian process model to equal the rolling of the mean of the Gaussian process in $\mathbb{R}^{d}$; (iii) to prescribe a finite-dimensional parameterisation of the rolled Gaussian process model, with statistical estimators and tests for the model parameters, in the practically useful setting when sample curves are observed at a finite set of times; and (iv) to establish conditions under which estimators of the parameters defining the mean curve consistently estimate their population counterpart. The technical contributions in (ii) concerning the interplay between notions of symmetry of the manifold and a distribution, and in (iv) on convergence of parallel transport maps along a sequence of random curves to their counterparts along the limiting curve, may be of independent interest.
@@ 2 Preliminaries
We record relevant definitions from Riemannian geometry and geometric statistics, and refer to Kobayashi and Nomizu (1996); Lee (2018) and Pennec et al. (2019), respectively, for details.
@@
Let $M$ be a complete connected Riemannian manifold of dimension $d$. Denote by $T_{p}M$ the tangent space at each point $p\in M$ and by $TM=\{T_{p}M:p\in M\}$ the collection of tangent spaces, known as the tangent bundle. On the tangent space $T_{p}M$ at each point $p$, we can define an inner product $\langle\cdot,\cdot\rangle_{p}:T_{p}M\times T_{p}M\rightarrow\mathbb{R}_{\geq 0}$ with induced norm $\|\cdot\|_{p}$; the collection $\{\langle\cdot,\cdot\rangle_{p}:p\in M\}$ is referred to as a Riemannian metric. The Riemannian metric varies smoothly along the manifold and allows us to measure distances, volumes, and angles.
@@
For a curve $\gamma:[0,1]\rightarrow M$, its velocity at $t$ is $\dot{\gamma}(t):=\mathrm{d}\gamma(t)/\mathrm{d}t\in T_{\gamma(t)}M$, and its length $L(\gamma)$ is the scalar $\int_{0}^{1}\|\dot{\gamma}(t)\|^{1/2}_{\gamma(t)}\text{d}t$. For all curves in this paper, the parameter domain is the unit interval, $[0,1]$. Often, but not always, $t$ is interpreted as time; we refer to it thus in the following.
@@
The Riemannian metric $g$ defines a unique affine connection called the Levi-Civita connection $\nabla$, known as the covariant derivative, which describes how vectors in tangent spaces change as one moves along a curve $\gamma$ from one point to another. A vector field $t\mapsto V(t)\in T_{\gamma(t)}M$ along a curve $\gamma$ is said to be parallel if $\nabla_{\dot{\gamma}(t)}V(t)=0$ for all $t$. Given a curve $\gamma:[0,1]\to M$ and $u\in T_{\gamma(t_{0})}M$ for $t_{0}\in[0,1]$, there exists a unique parallel vector field $V$ such that $V(t_{0})=u$.
@@
Vectors in different tangent spaces are identified using parallel transport along a curve $\gamma$: the parallel transport of $u\in T_{\gamma(s)}M$ along $\gamma$ to $T_{\gamma(t)}M$ is $V(t)$, where $V$ is the unique parallel vector field determined by $V(s)=u$. This engenders a linear isometry $P_{t\leftarrow s}^{\gamma}:T_{\gamma(s)}M\to T_{\gamma(t)}M,\quad P_{t\leftarrow s}^{\gamma}(u)=V(t)$, between tangent spaces at points along $\gamma$, such that $\langle u,w\rangle_{\gamma(s)}=\langle P_{t\leftarrow s}^{\gamma}(u),P_{t\leftarrow s}^{\gamma}(w)\rangle_{\gamma(t)}$ for every $u,w$; its inverse is denoted by $P_{s\leftarrow t}^{\gamma}$.
@@
Given a point $p$ and a geodesic $\gamma$ with $\gamma(0)=p$, a cut point of $p$ is defined as the point $\gamma(t_{0})$ such that $\gamma$ is a minimal geodesic on the interval $[0,t_{0}]$ but fails to be for $t>t_{0}$. The set of cut points of geodesics starting at $p$ is its cut locus $\mathcal{C}(p)$. The injectivity radius of $p$ is its distance to $\mathcal{C}(p)$, and the injectivity radius $i(M)$ of $M$ is the infimum of the injectivity radii of points in $M$.
The next two tools are necessary for moving between $M$ and its tangent spaces. Geodesics on $M$ are used to define the exponential map $\exp:TM\to M$, which maps a point $(p,v)$ in the tangent bundle to a point on $M$; in particular, when restricted to $T_{p}M$ it is defined via a geodesic $\gamma:[0,1]\to M$ starting at $\gamma(0)=p$ with initial velocity $\dot{\gamma}(0)=v$ as $\exp_{p}:T_{p}M\to M,\exp_{p}(v)=\gamma(1)$. From the Hopf-Rinow theorem, on a complete manifold the exponential map is surjective. On an star shape region around the origin in $T_{p}M$ it is a diffeomorphism onto its image outside of the cut locus of $p$, and a well-defined inverse $\exp^{-1}_{p}:M\to T_{p}M$ exists, known as the inverse exponential or logarithm map, that maps a point on $M$ outside of $\mathcal{C}(p)$ to $T_{p}M$; thus for any $q$ outside of $\mathcal{C}(p)$, $\rho(p,q)=\|\exp_{p}^{-1}(q)\|_{p}$. Relatedly, the tangent cut locus $\mathcal{T}(p)\subset T_{p}M$ of $p\in M$, is defined by the relation $\exp_{p}(\mathcal{T}(p))=\mathcal{C}(p)$.
There are many notions of curvature of a Riemannian manifold. We will mainly be concerned with sectional curvature at a point $p$, defined to be the Gaussian curvature at $p$ of the two-dimensional surface swept out by the set of all geodesics starting at $p$ with initial velocities lying in the two-dimensional subspace of $T_{p}M$ spanned by two linear independent vectors. Throughout, we will simply use curvature to mean sectional curvature, and describe a manifold $M$ as positively (non-positively) curved if its curvature at every point is positive (non-positive). Non-positively curved $M$ have infinite $i(M)$ with empty cut locus such that $\exp_{p}:T_{p}M\to M$ is a global diffeomorphism at every $p$.
The Fréchet mean of a probability measure $\nu$ on $M$ is defined as the minimizer of
@@
sufficient conditions that ensure existence and uniqueness of the Fréchet mean on manifolds are well-known (e.g. Afsari, 2011). When $\nu$ is the empirical measure of a sample of points in $M$, the minimizer is the sample Fréchet mean. Under suitable conditions the sample Fréchet mean converges almost surely to its population counterpart (Ziezold, 1977). Abusing terminology, we also speak of the Fréchet mean of a random variable $x$ where $x\sim\nu$; context will disambiguate the two. For a random curve $x:[0,1]\to M$ the population Fréchet mean curve, $\gamma_{\text{Fre}}:[0,1]\to M$, is defined as the Fréchet mean of $x$ pointwise in $t$; that is,
$\gamma_{\text{Fre}}(t)=\underset{s\in M}{\mathrm{argmin}}\thickspace E\{\rho^{2}(x(t),s)\}.$ (2.2)
For a sample of curves, $\{x_{1},\ldots x_{n}\}$, the sample Fréchet mean curve is $\hat{\gamma}_{\text{Fre}}:[0,1]\to M$,
$\hat{\gamma}_{\text{Fre}}(t):=\underset{s\in M}{\mathrm{argmin}}\frac{1}{n}\sum_{i=1}^{n}\rho^{2}(x_{i}(t),s).$ (2.3)
We will refer to a special class of manifolds known as symmetric manifolds when describing the relationship between the pointwise Fréchet mean curve and a mean curve under our proposed model. A symmetric manifold is a manifold $M$ in which for every $p\in M$ there is a unique isometry $\sigma_{p}:M\to M$ such that on $T_{p}M$ its differential, $\text{d}\sigma_{p}=-\text{id}$, the negative of the identity map; thus, $\sigma_{p}$ reverses geodesics $\gamma$ through $p$ so that if $\gamma(0)=p$, then $\sigma_{p}(\gamma(t))=\gamma(-t)$. A function $g:M\to\mathbb{R}$ on a symmetric manifold $M$ is said to be symmetric if $g(p)=g(\sigma_{p}(p))$ for every $p\in M$. Relatedly, a distribution $\nu$ on $M$ is said to possess even symmetry about $p\in M$, when $\nu=(\exp_{p})_{\#}\lambda$, the pushforward under the exponential map at $p$ of a mean-zero distribution $\lambda$ on $T_{p}M$ with Lebesgue density $f$, when $T_{p}M$ is identified with $\mathbb{R}^{d}$, satisfies $f(v)=f(-v)$ for every $v\in T_{p}M$. More generally, the measure $\nu$ on $M$ defined as a pushforward under the exponential map of a measure on $T_{p}M$ is commonly referred to as a wrapped distribution on $M$.
3 Rolled models for curves on $M$
3.1 Four maps for curves: unrolling, rolling, unwrapping, and wrapping

### v2

@@ Abstract
Given a planar curve, imagine rolling a sphere along that curve without slipping or twisting, and by this means tracing out a curve on the sphere. It is well known that such a rolling operation induces a local isometry between the sphere and the plane so that the two curves uniquely determine each other, and moreover, the operation extends to a general class of manifolds in any dimension. We use rolling to construct an analogue of a Gaussian process on a manifold starting from a Euclidean Gaussian process with mean $m$ and covariance $K$, and refer to it as a rolled Gaussian process parameterized by $m$ and $K$. The resulting model is generative, and is amenable to statistical inference given data as curves on a manifold. We identify conditions on the manifold under which the rolling of $m$ equals the Fréchet mean of the rolled Gaussian process, propose computationally simple estimators of $m$ and $K$, and derive their rates of convergence. We illustrate with examples on the unit sphere, symmetric positive-definite matrices, and with a robotics application involving 3D orientations.
@@ 1 Introduction
Many modern applications generate data where the statistical unit is a curve $t\mapsto x(t)\in M$ on a manifold $M$. Examples include data as trajectories on spheres (Su et al., 2014; Zhang et al., 2018a), shape spaces (Chevallier et al., 2021), the space of covariance matrices (Zhang et al., 2018b), and the space of probability densities (Chen et al., 2017). Despite prevalence of such data, there is a dearth of simple generative statistical models that are suitable, and amenable for inference, for discretely observed data curves.
@@
The four maps are compatible with the intrinsic geometry of $M$, and are optimal in a sense clarified below in not introducing distortions. This is in contrast to identifying $\mathbb{R}^{d}$ with a tangent space at a single fixed point on $M$, and using tangent space projections between the tangent space and $M$, and vice versa, which causes distortions. Our maps provide a natural way to define a covariance, and hence a Gaussian process-type model, on $M$, starting with one in $\mathbb{R}^{d}$.
1.1 Related work and contributions
@@
“Rolling” of a curve in $\mathbb{R}^{d}$ onto $M$, and its reverse operation, “unrolling”, originate from mathematical geometry, where they are known as antidevelopment and development respectively; see Kobayashi and Nomizu (1996) and Sharpe (2000). Jupp and Kent (1987) first exploited the idea in statistics, to define smoothing curves for time-indexed data points on the unit sphere. Kume et al. (2007) adapted their technique to shape spaces, and later Kim et al. (2021) generalized to Riemannian manifolds. These works, however, did not develop generative statistical models.
@@
Stochastic development and the definition of an intrinsic Brownian motion on $M$, (e.g. Hsu, 2002) and stochastic differential equations in $M$, can be used to define generative diffusion models for curves on $M$ (e.g. Sommer, 2019, 2020). However, the Markovian property of the models does not allow for rich covariance structures, and computational challenges abound: exact simulation is infeasible and numerical approximations are costly; likelihood functions are difficult to compute; and the infinitesimal formulation is not well-suited to modelling discretely observed curves.
Outside of the stochastic-differential-equation-based models, the recent papers have broadly focused on two different aspects of statistical analysis of data curves on $M$: (a) descriptive, where the focus has been on development and theoretical analysis of principal component analysis (PCA) with various notions of covariance of a random $M$-valued curve $x$ along a base curve $\gamma$, usually the Fréchet mean curve; (b) model-based, by wrapping a Gaussian vector field along a $\gamma$, but for which the practical and theoretical properties of estimators of model parameters are unclear. Relevant works with focus on aspect (a) include the extrinsic approach on embedded submanifolds $M$ by Dai and Müller (2019), and intrinsic approaches by Lin and Yao (2019) and Shao et al. (2022) with covariance defined using the family of tangent spaces along $\gamma$. Generality of the setting needed to develop the asymptotic theory of PCA in these papers impose restrictive assumptions that are usually difficult to verify in practice; for example, support of the distribution of $x$ is effectively restricted, either explicitly or via a condition on the Fréchet functional for $x$. In this setting, see also (Dubey and Müller, 2020, 2022) for curve methodology in general metric space, not explicity using the differentiable structure of a manifold, $M$. Works related to (b) include papers by Mallasto and Feragen (2018) and Wang et al. (2024) which either pre-specify $\gamma$ or estimate it separately from the model, and Liu et al. (2024) which is focused on the regression setting in the presence of Euclidean covariates. In these, theoretical support for the estimators is limited and convergence rates are absent.
Common to all papers that have focused on aspect (a) or (b) is a random vector field $V^{\gamma}$ along $\gamma$ with covariance defined with respect to the family of tangent spaces along $\gamma$. Reconciling such abstract definitions of covariance (e.g. Shao et al., 2022; Lin and Yao, 2019) with the practicality of needing coordinates for computation is a key complication when developing a unified modelling framework with theoretical guarantees.
The distinguishing feature of our approach lies in the joint specification of the base curve $\gamma$ and the vector field $t\mapsto V^{\gamma}(t)$ along $\gamma$ starting from a Euclidean Gaussian process in a single tangent space. The rolling procedure results in a $\gamma$ and a covariant Gaussian vector field $V^{\gamma}$ such that $V^{\gamma}(s)$ and $V^{\gamma}(t)$ are related by parallel transport along $\gamma$. Accordingly, our main contributions are as follows.
  • $\bullet$
    We use rolling to define an analogue of a Gaussian process $x$ on $M$ defined via the rolling of a Gaussian process $z$ with mean $m$ and covariance $K$ from a single tangent space, that is compatible with the intrinsic geometry of $M$. The model is equivariant to the choice of tangent space coordinates (Proposition 4.2).
  • $\bullet$
    We determine conditions on $M$ that ensure that the rolling of $m$ equals the Fréchet mean of $x$ (Theorem 4.1). From this, using the reverse operations of unrolling and unwrapping onto a single tangent space, we propose simple estimators of $m$ and $K$ without having to choose coordinates for the family of tangent spaces along a curve.
  • $\bullet$
    We derive rates of convergence of the estimators for fully observed curves (Theorem 5.2) that match the optimal rates in the Euclidean setting under assumptions that are easy to verify and satisfied by several manifolds of interest in statistics. We also derive two results that are new in the literature: (i) Aided by the properties of the Gaussian process $z$, we prove $C^{1}$ convergence of the sample Fréchet mean curve to its population counterpart (Theorem 5.1), which improves upon the present state-of-the-art uniform convergence (Dai and Müller, 2019; Lin and Yao, 2019); (ii) a holonomy bound with explicit cuvature-dependent constants for triangles in $M$, which is then used in deriving convergence rates for the estimators (Lemma C.3).
  • $\bullet$
    We consider a simple parametric rolled Gaussian model defined using basis functions in §4.1 that is tailored for discretely observed curves, and in §9 employ it on data curves of three-dimensional rotations arising from a novel robotics application.
Appendices A and B contain relevant definitions and computational details. Technical lemmas and proofs of all results are presented in Appendices C and D.
@@ 2 Preliminaries
We record relevant definitions from Riemannian geometry and refer to standard references (e.g Petersen, 2016; Lee, 2018) for details.
@@
Let $M$ be a complete connected Riemannian manifold of dimension $d$. Denote by $T_{p}M$ the tangent space at each point $p\in M$ and by $TM=\{T_{p}M:p\in M\}$ the collection of tangent spaces, known as the tangent bundle. On the tangent space $T_{p}M$ at each point $p$, we can define an inner product $\langle\cdot,\cdot\rangle_{p}:T_{p}M\times T_{p}M\rightarrow\mathbb{R}_{\geq 0}$ with induced norm $\|\cdot\|_{p}$; the collection $\{\langle\cdot,\cdot\rangle_{p}:p\in M\}$ is referred to as a Riemannian metric. The Riemannian metric varies smoothly along the manifold and allows us to measure distances, volumes, and angles. The operator norm for maps $B:T_{p}M\to T_{p}M$ is defined as $\|B\|_{\text{op}}:=\sup\{\|Bv\|_{p}:v\in T_{p}M,\|v\|_{p}=1\}$.
@@
For a curve $\gamma:[0,1]\rightarrow M$, with image $\gamma([0,1])\subset M$, its velocity at $t$ is $\dot{\gamma}(t):=\mathrm{d}\gamma(t)/\penalty 50\mathrm{d}t\in T_{\gamma(t)}M$, and its length $L(\gamma)$ is the scalar $\int_{0}^{1}\|\dot{\gamma}(t)\|^{1/2}_{\gamma(t)}\text{d}t$. For all curves in this paper, the parameter domain is the unit interval, $[0,1]$. Often, but not always, $t$ is interpreted as time; we refer to it thus in the following. Depending on the context, we will denote a curve by both $\gamma$ or $\gamma(t)$.
@@
Vector fields are smooth maps $X:M\to TM$ that map a point $p\in M$ to a vector $X(p)\in T_{p}M$. The Riemannian metric defines a unique affine connection called the Levi-Civita connection $\nabla$, known as the covariant derivative, which describes how vectors in tangent spaces change as one moves along a curve $\gamma$ from one point to another so that $\nabla_{X}Y$ is the covariant derivative of a vector field $Y$ with respect to another $X$. We use $\nabla_{p}X(p)\in T_{p}M$ to denote the covariant derivative of $\nabla_{X}X$ at the point $p$. Similarly, $\nabla_{p}^{2}X(p)$ is the second covariant derivative, or the Hessian, of the vector field $X$ at $p$.
@@
A vector field $t\mapsto V(t)\in T_{\gamma(t)}M$ along a curve $\gamma$ is said to be parallel if $\nabla_{\dot{\gamma}(t)}V(t)=0$ for all $t$. Given a curve $\gamma:[0,1]\to M$ and $u\in T_{\gamma(t_{0})}M$ for $t_{0}\in[0,1]$, there exists a unique parallel vector field $V$ such that $V(t_{0})=u$. Vectors in different tangent spaces are identified using parallel transport along a curve $\gamma$: the parallel transport of $u\in T_{\gamma(s)}M$ along $\gamma$ to $T_{\gamma(t)}M$ is $V(t)$, where $V$ is the unique parallel vector field determined by $V(s)=u$. This engenders a linear isometry $P_{s\rightarrow t}^{\gamma}:T_{\gamma(s)}M\to T_{\gamma(t)}M,\quad P_{t\leftarrow s}^{\gamma}(u)=V(t)$, between tangent spaces at points along $\gamma$, such that $\langle u,w\rangle_{\gamma(s)}=\langle P_{s\rightarrow t}^{\gamma}(u),P_{s\rightarrow t}^{\gamma}(w)\rangle_{\gamma(t)}$ for every $u,w\in T_{\gamma(s)}M$. Its inverse is denoted by $P_{s\leftarrow t}^{\gamma}$.
For a function $f:M\to\mathbb{R}$, we use $\nabla_{p}f(p)$ to denote the gradient of $f$ at $p$, which is defined implicity via the relation $\text{d}f(x)(v)=\langle\nabla_{p}f(p),v\rangle_{p}$, where $\text{d}f(x):T_{p}M\to\mathbb{R}$ is the directional derivative of $f$ at $p$ along $v$. The Hessian of $f$ at $p$ then is the bilinear map $\nabla^{2}_{p}f(p):T_{p}M\times T_{p}M\to\mathbb{R}$ with the corresponding Hessian operator $H(p):T_{p}M\to T_{p}M$. For $t\in[0,1]$, we abuse notation and also denote as $\nabla_{t}f(c(t))$ and $\nabla^{2}_{t}f(c(t))$, the gradient and Hessian of $f$ at $c(t)=p$, where $c:[0,1]\to M$ is a curve. The notation is helpful in representing derivatives of a function $f:M\times[0,1]\to\mathbb{R},(p,t)\mapsto f(p,t)$ with respect to either of the arguments.
@@
Given a point $p$ and a geodesic $\gamma$ with $\gamma(0)=p$, a cut point of $p$ is defined as the point $\gamma(t_{0})$ such that $\gamma$ is a minimal geodesic on the interval $0,t_{0}]$ but fails to be for $t>t_{0}$. The set of cut points of geodesics starting at $p$ is its cut locus $\mathcal{C}(p)$, and we also use $\mathcal{C}(A)=\{\mathcal{C}(p):p\in A\subseteq M\}$ to denote the the cut locus of a set. The injectivity radius $i(p)\in[0,\infty]$ of $p$ is its distance to $\mathcal{C}(p)$, and the injectivity radius $i(A)$ of $A\subseteq M$ is the infimum of the injectivity radii of points in $M$.
The next two tools are necessary for moving between $M$ and its tangent spaces. Geodesics on $M$ are used to define the exponential map $\exp:TM\to M$, which maps a point $(p,v)$ in the tangent bundle to a point on $M$; in particular, when restricted to $T_{p}M$ it is defined via a geodesic $\gamma:[0,1]\to M$ starting at $\gamma(0)=p$ with initial velocity $\dot{\gamma}(0)=v$ as $\exp_{p}:T_{p}M\to M,\exp_{p}(v)=\gamma(1)$. From the Hopf-Rinow theorem, on a complete manifold the exponential map is surjective. On an star shape region around the origin in $T_{p}M$ this map is a diffeomorphism onto its image outside of the cut locus of $p$, and a well-defined inverse $\exp^{-1}_{p}:M\to T_{p}M$ exists, known as the inverse exponential or logarithm map, that maps a point on $M$ outside of $\mathcal{C}(p)$ to $T_{p}M$; thus for any $q$ outside of $\mathcal{C}(p)$, $\rho(p,q)=\|\exp_{p}^{-1}(q)\|_{p}$. Relatedly, the tangent cut locus $\mathcal{T}(p)\subset T_{p}M$ of $p\in M$, is defined by the relation $\exp_{p}(\mathcal{T}(p))=\mathcal{C}(p)$.
There are many notions of curvature of a Riemannian manifold. We will mainly be concerned with sectional curvature at a point $p$, $\sec{(p)}$, defined to be the Gaussian curvature at $p$ of the two-dimensional surface swept out by the set of all geodesics starting at $p$ with initial velocities lying in the two-dimensional subspace of $T_{p}M$ spanned by two linear independent vectors. For $A\subset M$, the notation $\sec{(A)}$ represents $\{\sec(p):p\in M\}$ so that $k_{0}\leq\sec{(A)}\leq k_{1}$ implies that the sectional curvature of every $p\in A$ is upper and lower bounded by $k_{1}$ and $k_{0}$, respectively. Throughout, we will simply use curvature to mean sectional curvature, and describe a manifold $M$ as positively (non-positively) curved if its curvature at every point is positive (non-positive). Non-positively curved $M$ have empty cut loci with $i(M)=\infty$ so that $\exp_{p}:T_{p}M\to M$ is a global diffeomorphism at every $p$.
The Fréchet mean of a probability measure $\nu$ on $M$ is defined as the minimizer of Fréchet functional
@@
sufficient conditions that ensure existence and uniqueness of the Fréchet mean on manifolds are well-known (e.g. Afsari, 2011) and are assumed henceforth. When $\nu$ is the empirical measure of a sample of points in $M$, the minimizer of the Fréchet functional with respect to the empirical measure is the sample Fréchet mean. Under suitable conditions the sample Fréchet mean converges almost surely to its population counterpart (Ziezold, 1977). Abusing terminology, we also speak of the Fréchet mean of a random variable $x$ where $x\sim\nu$; context will disambiguate the two. For a random curve $x:[0,1]\to M$ the population Fréchet mean curve, $\gamma_{\text{Fre}}:[0,1]\to M$, is defined as the Fréchet mean of $x$ pointwise in $t$; that is, minimizing
$p\mapsto F(p,t):=\mathbb{E}[\rho^{2}(p,x(t))],$ (2.2)
For a sample of curves, $\{x_{1},\ldots x_{n}\}$, the sample Fréchet mean curve is $\hat{\gamma}_{\text{Fre}}:[0,1]\to M$, minimizing
$p\mapsto F_{n}(p,t):=\frac{1}{n}\sum_{i=1}^{n}\rho^{2}(p,x_{i}(t)).$ (2.3)
We assume that $\gamma_{\text{Fre}}$ exists and is unique for the rolled Gaussian model we propose (assumption A1). No such assumption will be made about its sample counterpart.
Two types of manifolds are of particular interest in this work, especially when deriving the asymptotic properties of estimators of the model parameters: Hadamard manifolds and symmetric spaces. Simply connected, complete manifolds with non-positive sectional curvatures everywhere are known as Hadamard manifolds. A symmetric space is a homogeneous manifold $M$ in which for every $p\in M$ there is a unique isometry $\sigma_{p}:M\to M$ such that on $T_{p}M$ its differential, $\text{d}\sigma_{p}=-\text{id}$, the negative of the identity map; thus, $\sigma_{p}$ reverses geodesics $\gamma$ through $p$ so that if $\gamma(0)=p$, then $\sigma_{p}(\gamma(t))=\gamma(-t)$. Symmetric spaces are classified as being of compact type or non-compact type, wherein the former is positively curved while the latter is non-positively curved. A function $g:M\to\mathbb{R}$ on a symmetric manifold $M$ is said to be symmetric if $g(p)=g(\sigma_{p}(p))$ for every $p\in M$. Relatedly, a distribution $\nu$ on $M$ is said to possess geodesic symmetry about $p\in M$, when $\nu=(\exp_{p})_{\#}\lambda$, where $\lambda$ is a mean-zero distribution on $T_{p}M$ with Lebesgue density $f$ that satisfies f(v)=f(v)f(v)=f(-v) for every $v\in T_{p}M$. The measure $\nu$ on $M$ defined as a pushforward under $\exp_{p}$ of a measure on $T_{p}M$ is commonly referred to as a wrapped distribution on $M$.
3 The unrolling, rolling, unwrapping, and wrapping maps